OpenStudy (anonymous):

The half-life of 234U, uranium-234, is 2.52 multiplied by 105 yr. If 97.6% of the uranium in the original sample is present, what length of time (to the nearest thousand years) has elapsed?

OpenStudy (amistre64):

\[A = Pe^{rt}\]\[A/P = e^{rt}\]\[ln(A/P) = rt\]\[\frac{ln(A/P)}{r} =t\]

OpenStudy (amistre64):

so it starts out as \[97.6=100e^{rt}\]